OpenStudy (anonymous):

What is the 7th term of the geometric sequence where a1 = -3 and a5 = -768?

OpenStudy (anonymous):

is it -3,072?

OpenStudy (helder_edwin):

if it is a geometruc sequence then since a1=-3 we have \(-768=a_5=a_1\cdot r^4\)

OpenStudy (helder_edwin):

u don't knw the ratio of the sequence but can find it using this equation

OpenStudy (cwrw238):

nth term = a1 * r^(n-1) r ^4 = a5 / a1 = -768 /- 3 = 256 so r = 4th root of 256

OpenStudy (anonymous):

After solving you get r = 4..

OpenStudy (cwrw238):

yea

OpenStudy (anonymous):

So: \[\large a_7 = (-3)(4)^{6} = ??\]